# Quantum entanglement transfer assisted via Duffing nonlinearity

### v2

@@ IV CONCLUSION
In this work, we investigated stationary entanglement in an optomechanical system involving a single optical field driving two coupled mechanical resonators. The inter-resonator phonon hopping rate was modulated to induce a synthetic magnetism effect. Additionally, one of the resonators incorporated a Duffing nonlinearity. Utilizing experimentally achievable parameters, we demonstrated a slight improvement in system stability, a favorable condition for enhancing stationary entanglement. For the case without the Duffing nonlinearity, the logarithmic negativity exhibited the phenomenon of sudden death and revival of entanglement. The entanglement reached optimal values near the synthetic phase of $\theta=m\frac{\pi}{2}$ (where $m$ is an odd integer) and vanished around $\theta=n\pi$ (where $n$ is any integer). Following a strategic selection of the synthetic phase, we observed a weakening and eventual vanishing of entanglement between the optical field and the Duffing resonator when the nonlinearity became sufficiently strong. However, there was a corresponding enhancement of entanglement with the nonlinearity-free resonator. This behavior suggests a process of entanglement transfer facilitated by the Duffing nonlinearity. Notably, the transferred entanglement displayed superior robustness against thermal fluctuations. Our findings pave the way for further exploration of entanglement transfer phenomena in complex systems. These results hold promise for generating robust and stable entanglement in nonlinear systems, which could prove instrumental for advancements in quantum information processing and quantum computing applications.
@@ Acknowledgments
This work has been carried out under the Iso-Lomso Fellowship at Stellenbosch Institute for Advanced Study (STIAS), Wallenberg Research Centre at Stellenbosch University, Stellenbosch 7600, South Africa. P. Djorwe acknowledges the receipt of a grant from the APS-EPS-FECS-ICTP Travel Award Fellowship Programme (ATAP), Trieste, Italy. A.K. Sarma acknowledges the STARS scheme, MoE, government of India (Proposal ID 2023-0161).

### v3

@@ IV CONCLUSION
In this work, we investigated stationary entanglement in an optomechanical system involving a single optical field driving two coupled mechanical resonators. The inter-resonator phonon hopping rate was modulated to induce a synthetic magnetism effect. Additionally, one of the resonators incorporated a Duffing nonlinearity. Utilizing experimentally achievable parameters, we demonstrated a slight improvement in system stability, a favorable condition for enhancing stationary entanglement. For the case without the Duffing nonlinearity, the logarithmic negativity exhibited the phenomenon of sudden death and revival of entanglement. The entanglement reached optimal values near the synthetic phase of $\theta=m\frac{\pi}{2}$ (where $m$ is an odd integer) and vanished around $\theta=n\pi$ (where $n$ is any integer). Following a strategic selection of the synthetic phase, we observed a weakening and eventual vanishing of entanglement between the optical field and the Duffing resonator when the nonlinearity became sufficiently strong. However, there was a corresponding enhancement of entanglement with the nonlinearity-free resonator. This behavior suggests a process of entanglement generation facilitated by the Duffing nonlinearity. Notably, the induced entanglement displayed superior resilience against thermal fluctuations. Our findings pave the way for further exploration of entanglement in complex systems. These results hold promise for generating strong and stable entanglement in nonlinear systems, which could prove instrumental for advancements in quantum information processing and quantum computing applications.
Appendix A Discussions on the experimental realization of our proposal
Our proposal can be realized both in optomechanical and electromechanical platforms. For optomechanical structures, our model can be thought as a photonic-crystal optomechanical circuit where both photonic and phononic hopping coupling are intrinsic to the system. Such a configuration has been recently experimentally implemented in Fang et al. (2017), where a reduced model consists of two optomechanical cavities in which the involved optical (mechanical) resonators are optically (mechanically) coupled through a photonic (phononic) hopping coupling. Under a large-detuning regime, one of the optical mode has been adiabatically eliminated, reducing the four mode system to a three-mode loop-coupled optomechanical system consisting of two mechanically coupled mechanical modes which are coupled to a single cavity-field mode as depicted in Figure 1c of our main text. Moreover, a driving of the optomechanical cavities with phase-correlated fields results to an effective synthetic magnetic field into the system. Similar procedure can be also engineered in a membrane-in-the-middle optomechanical systems, where two mechanically coupled mechanical resonators are located inside an optical cavity as the one shown in Figure 1b of our main text.
Figure 6: Investigation for nondegenerated mechanical resonators. (a) Entanglement between the electromagnetic field and the nonlinear resonator (mode $b_{1}$) versus the synthetic phase $\theta$. (b) Entanglement between the electromagnetic field and the linear resonator (mode $b_{2}$) versus the synthetic phase $\theta$. The solid line is without Duffing nonlinearity ($\eta=0$), while the dash-dotted and dashed lines are for $\eta=5\times 10^{-6}\omega_{m}$ and $\eta=5\times 10^{-4}\omega_{m}$, respectively. We have used $\kappa=2\times 10^{-1}\omega_{m}$, $J_{m}=2\times 10^{-1}\omega_{m}$, $\alpha^{in}=1000\omega_{m}^{1/2}$ and $n_{th}=100$. The mechanical frequencies and dissipations are $\omega_{1}=\omega_{m}$, $\gamma_{1}=\gamma_{m}$, $\omega_{2}=\frac{4\omega_{m}}{3}$, and $\gamma_{2}=\frac{4\gamma_{m}}{3}$. The other parameters are as in Fig.2.
Our proposed model can be also implemented in an electromechanical system as the one depicted in Figure 1a of our main text, which consists of multiple mechanical resonators driven by a common microwave cavity represented by an equivalent inductance L and capacitance C. The mechanical resonators involved are the moving capacitor gates, which are grounded to the substrate that is seen as the phononic hopping coupling. In such a system, the phase-dependent phonon-hopping interaction (i.e., the phase in a loop coupling induces an effective synthetic magnetism) between the nearest-neighboring mechanical resonators can be engineered by coupling them to a superconducting charge qubit. More details and calculations on this procedure can be found in Lai et al. (2022) of our manuscript.
It has been assumed in our investigation that the mechanical resonators involved are degenerated, that was to easy both our notations and calculations. However, we would like to mention that this assumption does not limit/destroy our results and observed phenomena. For that purpose, we have considered nondegenerated mechanical resonators and have displayed in Figure 6 a similar observed behaviors as depicted in Figure 3 of the main text. It can be seen that the observed phenomena are the same regardless the physical properties of the involved mechanical resonators. Such a conclusion reveals the fact that our investigation is not parameter-limitted, and can be flexibly reproduced experimentally.
@@ Acknowledgments
This work has been carried out under the Iso-Lomso Fellowship at Stellenbosch Institute for Advanced Study (STIAS), Wallenberg Research Centre at Stellenbosch University, Stellenbosch 7600, South Africa. P. Djorwe acknowledges the receipt of a grant from the APS-EPS-FECS-ICTP Travel Award Fellowship Programme (ATAP), Trieste, Italy. A.K. Sarma acknowledges the STARS scheme, MoE, government of India (Proposal ID 2023-0161). A.-H. Abdel-Aty is thankful to the Deanship of Graduate Studies and Scientific Research at University of Bisha for supporting this work through the Fast-Track Research Support Program.